OpenStudy (ryanchick):

please help me with this 2+3-4 divided by4x3x5x7x1+0

OpenStudy (anonymous):

so (2+3+-4)/(4*3*5*7*1+0) right?

OpenStudy (ryanchick):

i have no clue im completely lost with this problem

OpenStudy (anonymous):

well what is 2+3-4?

OpenStudy (ryanchick):

1

OpenStudy (anonymous):

what is 4*3*5*7

OpenStudy (anonymous):

4*3*5*7=(4*3)*(5*7)=12*35=?

OpenStudy (ryanchick):

@Bill_Cipher

OpenStudy (anonymous):

so what is 12*35?

OpenStudy (ryanchick):

420

OpenStudy (anonymous):

so if I interpreted the question correctly you should have the answer 1/420

OpenStudy (ryanchick):

okay think @Bill_Cipher